# DenseSplat: Densifying Gaussian Splatting SLAM with Neural Radiance Prior

### v1

@@ IV-B Evaluation of Scene Densification
TABLE VI: The mesh evaluation of our method and the Gaussian SLAM baselines on the Replica dataset [65]. The results are averaged over 8 scenes.
@@
NeRF NICE-SLAM [46] 2.85 3.02 89.34
Co-SLAM [48] 2.10 2.08 93.44
ESLAM [47] 0.97 1.05 98.60
3DGS SplaTAM [18] 2.74 4.02 84.89
MonoGS [19] 3.16 4.45 81.52
Photo-SLAM [20] 2.53 3.75 85.67
@@
Figure 5 illustrates the novel-view synthesis of DenseSplat and Gaussian-based SLAM systems on the Replica dataset [65]. The obstructed views typical of indoor environments pose remarkable challenges for Gaussian SLAM systems. For instance, SplaTAM [18] and MonoGS [19] exhibit notable deficiencies with holes in their reconstructed Gaussian maps. Although Photo-SLAM [20] mitigates these gaps using a pyramid feature extraction strategy, it results in sparser map representations, which diminish rendering quality. In contrast, DenseSplat employs the robust NeRF prior for Gaussian initialization, effectively filling these gaps by interpolating unobserved geometry. Our method also ensures a more uniform distribution of Gaussian primitives across the surface of the objects using our point sampling strategy, leading to better geometric alignment. Furthermore, Figure 5 also compares the rendering results of our method at a dense keyframe interval of 4, typical of current Gaussian SLAM systems, with a much sparser interval of 20. Despite the map being slightly less populated, DenseSplat continues to demonstrate robust reconstruction and interpolation capabilities. This effectiveness in sparse conditions significantly enhances real-time processing and, more importantly, relaxes the stringent requirement for dense multi-view observations previously essential in Gaussian SLAM systems. Figure 6 provides a broader view of the Replica scenes, showcasing the superior map completeness of our method compared to baseline approaches. The latter often display significant gaps due to common occlusions encountered in indoor environments.
Figure 6: Visualization of hole-filling in novel-view synthesis for the scene Room_0, Office_2, and Office_3 from the Replica dataset [65]. Employing the NeRF prior, our method effectively interpolates the unobserved geometry obscured by obstacles in the room, contrasting with Gaussian-based approaches that exhibit significant gaps, drastically compromising the completeness of the reconstruction.
Moreover, to quantitatively evaluate reconstruction completeness, we transform Gaussian maps into meshes using TSDF-fusion [69] and compare them with ground-truth meshes. As demonstrated in Table VI, DenseSplat significantly outperforms recent Gaussian SLAM systems in accuracy and scene completion ratio, which often struggle with gaps that result in underrepresented scenes.
Failure Case of Interpolation When the uncovered regions from viewpoints become extensively large, our method still faces challenges in completely filling holes. As illustrated in Figure 7, in the case of Apartment_2 from the Replica dataset [65], the limitations become apparent when the camera trajectory does not sufficiently cover the scene, leaving nearly half of the room uncaptured. This situation makes scene interpolation particularly challenging. The NeRF-based model such as Co-SLAM [48] manages to fill some gaps but introduces erroneous surfaces with significant artifacts. Our method outperforms other Gaussian-based methods by a large margin, using sparse primitives to fill gaps; however, it still leaves certain areas underrepresented. In scenarios where unobserved geometry primarily results from insufficient camera coverage, inferring these unobserved regions aligns more with scene extrapolation. Our method faces certain limitations as the NeRF prior is less reliable for extrapolation, and addressing these challenges will be a key direction for future research.
Figure 7: Visualization of hole-filling failures in novel-view synthesis for the scene Apartment_2 from the Replica dataset [65]. The camera trajectory is highlighted in red on the ground-truth map. Our method and Gaussian-based approaches struggle to completely fill the notably large gaps. In contrast, NeRF-based systems like Co-SLAM [48] manage to fill these gaps to some extent, though the results are not sufficiently smooth.
@@ V Ablation Study
In this section, we conduct extensive ablation studies on each core component of our system in terms of rendering quality, scene accuracy, and system efficiency in Table VII. By taking the tradeoff of rendering and mapping efficiency, our method capitalizes on the benefits of high-quality rendering using the NeRF model.
TABLE VII: Ablation study on Room0 of the Replica dataset [65]. To enable Gaussian initialization w/o NeRF, we randomly populate primitives on the space for the subsequent mapping process.
@@
full implementation 37.11 23.99 8.08 37.52 6.68
@@
Figure 8: Ablation study on keyframe intervals, presenting the mapping FPS and rendering PSNR as functions of keyframe intervals on the Room0 of the Replica dataset [65].
Ablation of Sampling Ratio Although precise control over the number of Gaussian primitives is challenging due to the densification process [3], adjusting the sampling ratio of the NeRF model roughly manages the granularity of the scene representation. As shown in Figure 9, compared to our original setting, downsampling the ratio by a factor of 4 results in a sparser primitive distribution, with a slight reduction in rendering quality.
Figure 9: Ablation study on sampling ratio, showing rendered images of Room0 from the Replica dataset [65], with centers $\mu$ shaded in blue. (a) Sampling ratio of 150 (b) Downsampled by 4
@@
Figure 10: Ablation study on BA, showing the rendered scenes of Apartment1 from the Replica dataset [65]. (a) baseline method employing naive frame-to-frame tracking. (b) no loop closure and BA on the tracking system. (c) no BA loss defined in Equation 9 and Equation 10. (d) Full implementation of DenseSplat.
@@ VI VR/AR Applications
Neural dense SLAM is well-suited for VR/AR applications [16], where the essential synthesis of novel views enables seamless and dynamic perspectives that adapt to user movements. Despite the substantial advantages offered by Gaussian-based systems, such as producing high-fidelity maps and efficient real-time rendering capabilities, they often display significant holes and gaps when confronted with limited viewpoint supervision, which can severely impact the user experience. As illustrated in Figure 11, we place the LEGO object reconstructed from the NeRF Synthetic Dataset [1] into the dense Gaussian map generated by SLAM systems. In comparison to previous methods like MonoGS [19], our method remarkably improves scene completeness by interpolating missing geometry, thereby enhancing the user’s immersion and interaction within the virtual environment.
Figure 11: Visulization of the VR/AR demo on Office_3 of the Replica dataset [65]. We introduce the reconstructed LEGO object from the NeRF Synthetic Dataset [1] into the real-time reconstructed map and show the novel-view synthesis result.
@@ VII Conclusion
Limitation DenseSplat also has its limitations. As discussed in Section IV-B, its scene interpolation capabilities are dependent on the NeRF model, and thus inherit NeRF’s limitations. When the missing areas become extensively large and the NeRF model cannot adequately capture the geometry, both our method and NeRF struggle to fill gaps, leaving some regions underrepresented on the reconstructed map. Additionally, because DenseSplat employs explicit Gaussian primitives for scene representation, storing the high-fidelity map requires more memory compared to NeRF models, which use neural implicit representations. Moreover, while DenseSplat supports submap systems to minimize memory usage during computation, there is room for a more advanced submap strategy that could potentially be expanded in real-world multi-agent systems. Addressing these limitations will be the focus of future research.

### v2

@@ IV-B Evaluation of Scene Densification
Figure 6 illustrates the novel-view synthesis of DenseSplat and Gaussian-based SLAM systems on the Replica dataset [straub2019replica]. The obstructed views typical of indoor environments pose remarkable challenges for Gaussian SLAM systems. For instance, SplaTAM [keetha2024splatam] and MonoGS [matsuki2024gaussian] exhibit notable deficiencies with holes in their reconstructed Gaussian maps. Although Photo-SLAM [huang2024photo] mitigates these gaps using a pyramid feature extraction strategy, it results in sparser map representations, which diminish rendering quality. In contrast, DenseSplat employs the robust NeRF prior for Gaussian initialization, effectively filling these gaps by interpolating unobserved geometry. Our method also ensures a more uniform distribution of Gaussian primitives across the surface of the objects using our point sampling strategy, leading to better geometric alignment. Furthermore, Figure 6 also compares the rendering results of our method at a dense keyframe interval of 4, typical of current Gaussian SLAM systems, with a much sparser interval of 20. Despite the map being slightly less populated, DenseSplat continues to demonstrate robust reconstruction and interpolation capabilities. This effectiveness in sparse conditions significantly enhances real-time processing and, more importantly, relaxes the stringent requirement for dense multi-view observations previously essential in Gaussian SLAM systems. Figure 7 provides a broader view of the Replica scenes, showcasing the superior map completeness of our method compared to baseline approaches. The latter often display significant gaps due to common occlusions encountered in indoor environments.
TABLE VI: The mesh evaluation of our method and the Gaussian SLAM baselines on the Replica dataset [straub2019replica]. The results are averaged over 8 scenes.
@@
NeRF NICE-SLAM [zhu2022nice] 2.85 3.02 89.34
Co-SLAM [wang2023co] 2.10 2.08 93.44
ESLAM [johari2023eslam] 0.97 1.05 98.60
3DGS SplaTAM [keetha2024splatam] 2.74 4.02 84.89
MonoGS [matsuki2024gaussian] 3.16 4.45 81.52
Photo-SLAM [huang2024photo] 2.53 3.75 85.67
@@
Moreover, to quantitatively evaluate the reconstruction completeness, we transform Gaussian maps into meshes using TSDF-fusion [curless1996volumetric] and compare them with ground-truth meshes. As demonstrated in Table VI, DenseSplat significantly outperforms recent Gaussian SLAM systems in accuracy and scene completion ratio, which often struggle with gaps that result in underrepresented scenes.
Figure 8: Visualization of the ablation study on Gaussian primitive initialization approaches for Room0 of the Replica dataset [straub2019replica]. For direct backprojection, we backproject the RGB-D frames and downsample by $\times 1$ to initialize primitives. For native NeRF-based sampling, we uniformly sample NeRF using a grid size of 0.01. For our geometry‐aware sampling, we use $\tau_{\rm grid}=0.001$ and the same sampling grid size for initialization. Our strategy achieves superior geometric accuracy compared to direct backprojection, which exhibits substantial holes, and the naive NeRF‐based method, which produces many redundant floating primitives.
@@ V Ablation Study
TABLE VII: Ablation study of each component of DenseSplat on Room0 of the Replica dataset [straub2019replica].
@@
Ours 37.11 23.99 8.08 37.52 6.68
In this section, we conduct extensive ablation studies on each hyperparameter and core component of our system in terms of rendering quality, scene accuracy, and system efficiency. The overall ablation analysis is presented in Table VII. By taking the tradeoff of rendering and mapping efficiency, our method capitalizes on the benefits of high-quality rendering using the NeRF model.
TABLE VIII: Ablation study on the initialization of primitives $\mathcal{G}$, conducted on Room0 from the Replica dataset [straub2019replica]. For the backprojection-based initialization methods shown in the middle rows, we directly backproject pixels from the RGB-D input frame, using downsampling ratios of $\times 1$, $\times 8$, $\times 16$.
[…]
Random initialized $\mathcal{G}$ 18.19 117.47 10.12 190.64 17.13
Backprojected $\mathcal{G}$ (down $\times 1$}) 36.26 39.96 8.33 512.04 23.49
Backprojected $\mathcal{G}$ (down $\times 8$}) 33.27 57.13 11.56 339.17 20.39
Backprojected $\mathcal{G}$ (down $\times 16$}) 31.69 92.17 15.77 205.82 18.96
Naive NeRF sampling 34.12 16.54 6.64 289.36 21.31
Geometry-aware sampling (ours) 37.11 23.99 8.08 37.52 6.68
Ablation of NeRF-based Initialization To evaluate the effectiveness of our geometry-aware sampling strategy, Table VIII presents an ablation study on different primitive initialization strategies. We pay particular attention to the method of direct backprojection from RGB-D streams, which is commonly applied in previous Gaussian-based SLAM systems [keetha2024splatam, matsuki2024gaussian, yugay2023gaussian]. Specifically, the baseline approaches include random primitive initialization, direct backprojection of RGB-D frames downsampled by $\times 1$, $\times 8$, and $\times 16$, and a naive NeRF-based sampling approach that uniformly samples NeRF without removing redundant points. The ablation results shows that, by actively selecting surface transition points, our geometry-aware point sampling strategy achieves optimal rendering quality while consuming the least memory.
Furthermore, Figure 8 shows novel-view synthesis results for Room0 using direct backprojection, naive NeRF-based sampling, and our adaptive strategy. Direct backprojection leaves substantial holes due to insufficient or occluded views, while naive NeRF-based sampling introduces floating primitives due to the lack of filtering. Our method delivers a complete reconstruction, effectively filling gaps and maintaining geometric accuracy.
Ablation of Thresholding Hyperparameters We investigate three thresholding hyperparameters in Table IX: $\tau_{\rm grid}$ for NeRF sampling, $\tau_{\rm prune}$ for primitive pruning, and $\tau_{\rm scale}$ for scale regularization. Setting $\tau_{\rm grid}=0$ is equivalent to naive NeRF sampling. We employ a small sampling threshold $\tau_{\rm grid}=0.001$ to effectively remove erroneous sampling points from the NeRF prior, improving geometry accuracy and system efficiency. The parameter $\tau_{\rm prune}$ controls the removal of primitives with negligible radiance contribution; a small threshold brings certain improvement in quality and efficiency, while a larger value remarkably accelerates rendering but degrades performance by pruning too many primitives. The parameter $\tau_{\rm scale}$ penalizes excessively large Gaussian primitives to suppress view-dependent artifacts, especially in sparse-view settings. Smaller thresholds encourage finer primitives at the cost of increased memory usage, whereas $\tau_{\rm scale}=1$ offers a balanced trade-off between primitive granularity and overall efficiency.
TABLE IX: Ablation study of thresholding hyperparameters, conducted on Room0 from the Replica dataset [straub2019replica]. Note that $\tau_{\rm grid}=0$ is equivalent to naive NeRF-based initialization.
[…]
$\tau_{\rm grid}=0$ 34.12 16.54 6.64 289.36 21.31
$\tau_{\rm grid}=0.001$ (ours) 37.11 23.99 8.08 37.52 6.68
$\tau_{\rm grid}=0.005$ 36.58 23.07 7.61 55.97 7.74
$\tau_{\rm grid}=0.01$ 35.56 20.15 6.70 78.63 9.87
$\tau_{\rm prune}=0$ 31.97 20.12 6.21 299.36 21.67
$\tau_{\rm prune}=0.001$ (ours) 37.11 23.99 8.08 37.52 6.68
$\tau_{\rm prune}=0.01$ 35.19 31.91 11.54 21.23 5.08
$\tau_{\rm scale}=0.01$ 35.15 17.77 5.34 160.77 11.25
$\tau_{\rm scale}=0.1$ 37.20 20.94 6.89 70.61 9.34
$\tau_{\rm scale}=1$ (ours) 37.11 23.99 8.08 37.52 6.68
Figure 9: Ablation study on sampling grid size, showing rendered images of Room0 from the Replica dataset [straub2019replica], with primitive’s centers $\mu$ shaded in blue. (a) Sampling size of 0.01 (b) Sampling size of 0.001.
TABLE X: Ablation study of sampling grid size during NeRF sampling, conducted on Room0 from the Replica dataset [straub2019replica].
[…]
$\text{grid\_size}=0.001$ 36.28 21.15 6.93 69.66 9.27
$\text{grid\_size}=0.01$ (ours) 37.11 23.99 8.08 37.52 6.68
$\text{grid\_size}=0.1$ 35.44 27.26 9.88 25.91 5.37
Ablation of Sampling Size Although precise control over the number of Gaussian primitives is difficult due to the densification process [kerbl20233d], adjusting the sampling grid size of the NeRF model can roughly control the granularity of the scene representation. We perform an ablation study on the sampling grid size, as shown in Table X. The corresponding visualization in Figure 9 shows that, compared to our choice of 1 cm, increasing the sampling grid size by a factor of 10 produces a sparser primitive distribution and reduces rendering quality to some extent.
@@
Figure 10: Ablation study on keyframe intervals, presenting the mapping FPS and rendering PSNR as functions of keyframe intervals on the Room0 of the Replica dataset [straub2019replica].
@@
Figure 11: Ablation study on BA, showing the rendered scenes of Apartment1 from the Replica dataset [straub2019replica]. (a) baseline method employing naive frame-to-frame tracking. (b) no loop closure and BA on the tracking system. (c) no BA loss defined in Equation 9 and Equation 10. (d) Full implementation of DenseSplat.
Figure 12: Visualization of hole-filling failures in novel-view synthesis for the scene Apartment_2 from the Replica dataset [straub2019replica]. The camera trajectory is highlighted in red on the ground-truth map. Our method and Gaussian-based approaches struggle to completely fill the notably large gaps. In contrast, NeRF-based systems like Co-SLAM [wang2023co] manage to fill these gaps to some extent, though the results are not sufficiently smooth.
Analysis of an Interpolation Failure Case When the uncovered regions from viewpoints become extensively large, our method still faces challenges in completely filling holes. As illustrated in Figure 12, in the case of Apartment_2 from the Replica dataset [straub2019replica], the limitations become apparent when the camera trajectory does not sufficiently cover the scene, leaving nearly half of the room uncaptured. This situation makes scene interpolation particularly challenging. The NeRF-based model, such as Co-SLAM [wang2023co], manages to fill some gaps but introduces erroneous surfaces with significant artifacts. Our method outperforms other Gaussian-based methods by a large margin, using sparse primitives to fill gaps; however, it still leaves certain areas underrepresented. In scenarios where unobserved geometry primarily results from insufficient camera coverage, inferring these unobserved regions aligns more with scene extrapolation. Our method faces certain limitations as the NeRF prior is less reliable for extrapolation, and addressing these challenges will be a key direction for future research.
@@ VI VR/AR Applications
Neural dense SLAM is well-suited for VR/AR applications [zhai2024nis], where the essential synthesis of novel views enables seamless and dynamic perspectives that adapt to user movements. Despite the substantial advantages offered by Gaussian-based systems, such as producing high-fidelity maps and efficient real-time rendering capabilities, they often display significant holes and gaps when confronted with limited viewpoint supervision, which can severely impact the user experience. As illustrated in Figure 13, we place the LEGO object reconstructed from the NeRF Synthetic Dataset [mildenhall2021nerf] into the dense Gaussian map generated by SLAM systems. In comparison to previous methods like MonoGS [matsuki2024gaussian], our method remarkably improves scene completeness by interpolating missing geometry, thereby enhancing the user’s immersion and interaction within the virtual environment.
Figure 13: Visulization of the VR/AR demo on Office_3 of the Replica dataset [straub2019replica]. We introduce the reconstructed LEGO object from the NeRF Synthetic Dataset [mildenhall2021nerf] into the real-time reconstructed map and show the novel-view synthesis result.
@@ VII Conclusion
Limitation DenseSplat also has its limitations. As discussed in Section IV-B, its scene interpolation capabilities are dependent on the NeRF model, and thus inherit NeRF’s limitations. When the missing areas become excessively large and the NeRF model cannot adequately capture the geometry, both our method and NeRF struggle to extrapolate the scene geometry, leaving some regions underrepresented on the reconstructed map. Although recent efforts using generative priors show potential, extrapolation under realistic and large-scale conditions remains an open challenge. Additionally, because DenseSplat employs explicit Gaussian primitives for scene representation, storing the high-fidelity map requires more memory compared to NeRF models, which use neural implicit representations. Moreover, while DenseSplat supports submap systems to minimize memory usage during computation, there is room for a more advanced submap strategy that could potentially be expanded in real-world multi-agent systems. Addressing these limitations will be the focus of future research.
[Uncaptioned image] Mingrui Li is currently pursuing the Ph.D. degree with the School of Information and Communication Engineering, Dalian University of Technology, Dalian, China. His research interests include 3D reconstruction, simultaneous localization and mapping (SLAM), and computer vision.
[Uncaptioned image] Shuhong Liu is currently pursuing the Ph.D. degree in Department of Mechano-informatics, Information Science and Technology with the University of Tokyo, Tokyo, Japan. Before that, he received his bachelor’s degree in Department of Electrical and Computer Engineering at University of Waterloo, Ontario, Canada, and the master’s degree in Creative Informatics, Information Science and Technology, the University of Tokyo, Tokyo, Japan. His research interests include 3D computer vision, visual SLAM, and computational photography.
[Uncaptioned image] Tianchen Deng is currently pursuing the Ph.D. degree in control science and engineering with Shanghai Jiao Tong University, Shanghai, China. His main research interests include 3D Reconstruction, long-term visual simultaneous localization and mapping (SLAM), and vision-based localization.
[Uncaptioned image] Hongyu Wang (Member, IEEE) received his B.S. degree in electronic engineering from the Jilin University of Technology, Changchun, China, in 1990, the M.S. degree in electronic engineering from the Graduate School, Chinese Academy of Sciences, Beijing, China, in 1993, and the Ph.D. degree in precision instrument and optoelectronics engineering from Tianjin University, Tianjin, China, in 1997. He is currently a Professor with the Dalian University of Technology, Dalian, China. His research interests include image processing, computer vision, 3-D reconstruction, and simultaneous localization and mapping (SLAM).